# The entries of the Sinkhorn limit of an $m\times n$ matrix

### v1

@@ 2. The Sinkhorn limit of a $3\times 3$ matrix
Begin with a general $3\times 3$ matrix $A$ with symbolic entries; in particular, there are no algebraic relations between the entries. Let $r,s$ be symbols; eventually we will set $r$ to be the scalar factor $\frac{a_{31}}{a_{21}}$. Apply Theorem 6 to $A$ to obtain a polynomial equation satisfied by the top left entry of $\operatorname{Sink}(A)$. Then replace each instance of $\det A$ in this polynomial with $s\det A_{\{2,3\},\{2,3\}}$. By the definition of $M(S)$, each monomial now contains $\det A_{\{2,3\},\{2,3\}}$ as factor; divide by this factor. Finally, replace $a_{3,j}$ with $ra_{2,j}$ for each $j\in\{1,2,3\}$. The resulting polynomial factors as a product two cubic polynomials in $x$. One of these cubic factors is independent of $r$ and $s$; this is the polynomial $e_{3}x^{3}+e_{2}x^{2}+e_{1}x+e_{0}$.
To justify the construction, let $A$ now be the matrix in the statement of the corollary, let $r=\frac{a_{31}}{a_{21}}$ be the scalar factor, and fix a real number $s\neq 0$. We approximate $A$ by matrices with no 00 minors. Namely, let $B(t)$ be a continuous ($3\times 3$ matrix)-valued function such that $\lim_{t\to 0^{+}}B(t)=A$ and, for all $t>0$, no minor of $B(t)$ is 00. Further, we assume for all $t>0$ that $\frac{\det B(t)}{\det B(t)_{\{2,3\},\{2,3\}}}=s$. Apply Theorem 6 to $B(t)$ where $t>0$. Since $\det B(t)=s\det B(t)_{\{2,3\},\{2,3\}}$ by assumption, the polynomial given by Theorem 6 is of the form $(\det B(t)_{\{2,3\},\{2,3\}})g(t,x)$. The entries of $\operatorname{Sink}(B(t))$ are continuous functions of the entries of $B(t)$, and the roots of a polynomial are continuous functions of its coefficients, so $g(x)\mathrel{\mathop{:}}=\lim_{t\to 0^{+}}g(t,x)$ is a polynomial for the top left entry of $\operatorname{Sink}(A)$. Moreover, since only one of the two cubic factors of $g(x)$ is independent of $s$, the top left entry of $\operatorname{Sink}(A)$ is a root of that cubic factor. ∎
@@ 3. Square matrices and solving for coefficients
In this section, we use Theorem 6 to infer the form of an equation satisfied by the top left entry of the Sinkhorn limit of a positive $n\times n$ matrix. We then interpolate the coefficients in the equation for $4\times 4$ matrices to obtain Conjecture 10.
For a positive $2\times 2$ matrix $A$, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies Equation (1), namely $\left(a_{11}a_{22}-a_{12}a_{21}\right)x^{2}-2a_{11}a_{22}\,x+a_{11}a_{22}=0$. Multiplying by $a_{11}$, this can be written using $\Delta$ and $\Gamma$ as
@@
Let $n\geq 1$. There exist integers $c_{S}(n)$, indexed by subsets $S\subseteq D(n,n)$, such that, for every positive $n\times n$ matrix $A$, the top left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
This conjecture predicts that, since both $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ are homogeneous degree-$(\lvert{R}\rvert+1)$ polynomials in the entries of $A$, each monomial $M(S)$ (and therefore also the coefficient of each $x^{k}$ in Equation (8)) is a homogeneous polynomial with degree
$\sum_{(R,C)\in D(n,n)}(\lvert{R}\rvert+1)=\sum_{i=0}^{n-1}\binom{n-1}{i}^{2}(i+1)=\frac{n+1}{2}\binom{2n-2}{n-1}.$
For $n=1,2,3,\dots$, this degree is $1,3,12,50,210,882,3696,15444,\dots$ [21, A092443].
We mention two surprising properties of the polynomial in Theorem 6 that we expect to generalize to the polynomial in Conjecture 9. The first is a symmetry. Each coefficient $d_{k}$ in Theorem 6 can be obtained from $d_{6-k}$ by replacing
@@
The second is that, for each $k$, the sum of the coefficients $c_{S}(3)$ for $\lvert{S}\rvert=k$ is a signed binomial coefficient, namely
$\sum_{S\subseteq D(3,3)}c_{S}(3)x^{\lvert{S}\rvert}=(x-1)^{6}.$
Analogous statements also hold for the quadratic polynomial in Equation (6); each coefficient $d_{k}$ is related to $d_{2-k}$ by a symmetry, and
$\sum_{S\subseteq D(2,2)}c_{S}(2)x^{\lvert{S}\rvert}=(x-1)^{2}.$
@@
It remains to determine the coefficients $c_{S}(n)$. They are not uniquely determined by the conditions in Conjecture 9 since we can scale the polynomial. To remove this source of non-uniqueness, we define $c_{\{\}}(n)=1$ for all $n\geq 1$. Based on the polynomials for $n=2$ and $n=3$, we conjecture that $c_{D(n,n)}(n)=1$ for all $n\geq 2$. For $n=3$ we determined the coefficients $c_{S}(3)$ from the output of a Gröbner basis computation, but for $n\geq 4$ this computation seems to be infeasible. For a general $4\times 4$ matrix, we aborted the computation after $1$ week.
Instead, we generate many pseudorandom $n\times n$ matrices $A$, identify the top left entry of $\operatorname{Sink}(A)$ as an algebraic number for each, and set up systems of linear equations in $c_{S}(n)$. We describe these three steps next.
@@
In the second step, for each matrix $A$ generated in the first step, we determine a polynomial equation satisfied by the top left entry $x$ of $\operatorname{Sink}(A)$. There are two possible methods. One method is to use Gröbner bases; this is faster than the Gröbner basis computation for a matrix with symbolic entries, but for $n\geq 5$ it is still slow. Therefore we use another method, which is to apply the iterative scaling process to obtain a numeric approximation to $\operatorname{Sink}(A)$ and then guess a polynomial for its top left entry. We begin by numericizing the integer entries of $A$ to high precision. For $n=4$ we use precision $2^{12}$, and for $n=5$ we use precision $2^{15}$. Then we iteratively scale until we reach a fixed point. The precision of the entries drops during the scaling process, but with the initial precisions $2^{12}$ and $2^{15}$ we get entries with sufficiently high precision that we can reliably recognize them using PSLQ. The expected degree of the polynomial is $d=\binom{2n-2}{n-1}$ according to Conjecture 9. Building in redundancy, we use Mathematica’s RootApproximant [23] to approximate $x$ by an algebraic number with target degree $d+2$. When the output has degree $d$, which is almost always the case, this is strong evidence that the approximation is in fact the exact algebraic number we seek. Occasionally the output has degree less than $d$, in which case we discard the matrix; for example, one $4\times 4$ matrix produced a polynomial with degree $8$ rather than $20$, presumably because the general degree-$20$ polynomial, when evaluated at the entries of the matrix, is reducible. Finally, we perform a check on the output by computing the ratio of its leading coefficient to its constant coefficient. This ratio should be $\frac{M(D(n,n))}{M(\{\})}$, assuming Conjecture 9 is correct and $c_{D(n,n)}(n)=1$. All outputs passed this test. We record the matrix $A$ along with the polynomial equation satisfied by $x$, and this will give us $1$ equation in the third step. (In fact we can get $n^{2}$ equations by applying PSLQ to each entry of the numeric approximation to $\operatorname{Sink}(A)$ and recording each polynomial along with the matrix obtained by swapping the appropriate rows and columns of $A$. For $n=5$ this is worthwhile, since our implementation of the iterative scaling process takes roughly $3$ minutes to reach a fixed point.)
@@
The third step is to determine the coefficients $c_{S}(n)$ in the coefficient of $x^{k}$ in Conjecture 9. For a given $k$, we do this by solving a system of linear equations involving the coefficients $c_{S}(n)$ where $\lvert{S}\rvert=k$. We assume $c_{T}(n)=c_{S}(n)$ if $T\equiv S$, so it suffices to determine $c_{S}(n)$ for one representative $S$ of each equivalence class, analogous to Theorem 6. This reduces the number of unknown coefficients, which reduces the number of equations we need, which reduces the number of matrices $A$ we apply the iterative scaling process to in the second step above. However, first we must partition $\{S\subseteq D(n,n):\lvert{S}\rvert=k\}$ into its equivalence classes under $\equiv$. Some care must be taken to do this efficiently; we make use of the fact that row permutations commute with column permutations, so it suffices to apply row permutations first. Once we have computed the equivalence classes, we take each polynomial computed in the second step above, scale it by an integer so that its leading coefficient is $M(D(n,n))$ (and therefore its constant coefficient is $M(\{\})$), extract the coefficient of $x^{k}$, and set this coefficient equal to $\sum_{S}c_{S}(n)\Sigma(S)$ where the sum is over one representative from each equivalence class of size-$k$ subsets. The number of equivalence classes tells us how many such equations we need in order to solve for the unknown coefficients $c_{S}(n)$. We include more equations than necessary in the system, building in redundancy, so that if a solution is found then we can be confident that the conjectured form is correct. Then we solve the system. In practice, even setting up the system can be computationally expensive. For $n=5$ and $k=4$, our initial implementation took $6$ days to set up the system (before solving!) because there are $1518$ unknown coefficients $c_{S}(n)$, so we need at least that many equations, and each equation involves $\binom{70}{4}=916895$ monomials, each of which is a product of $70$ determinants. For small $k$, a more efficient way to construct each equation is to take advantage of the fact that, for each pair $S,T$ of subsets of $D(5,5)$, the products $M(S)$ and $M(T)$ have most factors in common, and almost all are $\Gamma$ factors. Therefore, we form an equivalent but substantially simpler equation by dividing both sides by $M(\{\})=\prod_{(R,C)\in D(5,5)}\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$. On the left, we compute $M(\{\})$ once and divide the extracted coefficient of $x^{k}$ by $M(\{\})$. On the right, we divide each $M(S)$ by $M(\{\})$. Instead of computing $M(S)/M(\{\})$ from definitions, we precompute the ratio $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)/\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ for each $(R,C)\in D(5,5)$; then, for each $S$ with $\lvert{S}\rvert=k$, we use the precomputed ratios to compute
@@
Here we’re using the fact that no minor is 00 to divide by $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$. This method is also used by SinkhornPolynomials to compute polynomials more quickly when no minor is 00.
We now carry out these three steps for $n=4$ to obtain a polynomial for the top left entry $x$ of $\operatorname{Sink}(A)$ for $4\times 4$ matrices $A$. This polynomial is analogous to the polynomial in Theorem 6 for $3\times 3$ matrices. The number of equivalence classes of size-$k$ subsets of $D(4,4)$ for $k=0,1,\dots,20$ is
@@
The number of unknown coefficients $c_{S}(4)$ is the sum of these numbers, which is $17920$. We use the definition $c_{\{\}}(4)=1$ and the conjecture $c_{D(4,4)}(4)=1$, and we solve the remaining $19$ systems of linear equations, the largest of which requires $3024$ equations. Unfortunately, for each $k\in\{4,5,\dots,16\}$ the system has multiple solutions. For example, when we solve the system for $k=4$, only $104$ of the $123$ unknown coefficients $c_{S}(4)$ with $\lvert{S}\rvert=4$ are uniquely determined; the other $19$ are parameterized by $2$ free variables. For $k=10$, the solution space has dimension $1141$. This is not a weakness of the interpolation strategy but rather implies that the coefficients of $x^{4},x^{5},\dots,x^{16}$ have multiple representations as linear combinations of class sums $\Sigma(S)$ and therefore that relations exist among these class sums. However, by setting the free variables to 00 (or any other values), we obtain variable-free coefficients $c_{S}(4)$.
@@
determine a polynomial equation satisfied by the top left entry $x$ of the Sinkhorn limit of a $4\times 4$ matrix, namely
$\sum_{k=0}^{20}\left(\sum c_{S}(4)\Sigma(S)\right)x^{k}=0$
@@
and let $x$ be the top left entry of $\operatorname{Sink}(A)$. Conjecture 10 gives
@@
In principle, we can use the same method to interpolate a polynomial equation for $n\times n$ matrices for any given $n$. However, the computation is formidable. For $5\times 5$ matrices, we were able to compute the coefficients for $k=4$ and $k=5$ (where the families of representations are respectively $8$-dimensional and $44$-dimensional), but for $k=6$ we could not get the computation of equivalence classes to finish (aborting after $2$ weeks). Completing the computation up to the halfway point $k=\frac{1}{2}\binom{8}{4}=35$, after which we could use the conjectural symmetry, seems to be infeasible.

### v2

@@ 2. The Sinkhorn limit of a $3\times 3$ matrix
Begin with a general $3\times 3$ matrix $A$ with symbolic entries; in particular, there are no algebraic relations between the entries. Let $r,s$ be symbols; eventually we will set $r$ to be the scalar factor $\frac{a_{31}}{a_{21}}$. Apply Theorem 6 to $A$ to obtain a polynomial equation satisfied by the top-left entry of $\operatorname{Sink}(A)$. Then replace each instance of $\det A$ in this polynomial with $s\det A_{\{2,3\},\{2,3\}}$. By the definition of $M(S)$, each monomial now contains $\det A_{\{2,3\},\{2,3\}}$ as factor; divide by this factor. Finally, replace $a_{3,j}$ with $ra_{2,j}$ for each $j\in\{1,2,3\}$. The resulting polynomial factors as a product two cubic polynomials in $x$. One of these cubic factors is independent of $r$ and $s$; this is the polynomial $e_{3}x^{3}+e_{2}x^{2}+e_{1}x+e_{0}$.
To justify the construction, let $A$ now be the matrix in the statement of the corollary, let $r=\frac{a_{31}}{a_{21}}$ be the scalar factor, and fix a real number $s\neq 0$. We approximate $A$ by matrices with no 00 minors. Namely, let $B(t)$ be a continuous ($3\times 3$ matrix)-valued function such that $\lim_{t\to 0^{+}}B(t)=A$ and, for all $t>0$, no minor of $B(t)$ is 00. Further, we assume for all $t>0$ that $\frac{\det B(t)}{\det B(t)_{\{2,3\},\{2,3\}}}=s$. Apply Theorem 6 to $B(t)$ where $t>0$. Since $\det B(t)=s\det B(t)_{\{2,3\},\{2,3\}}$ by assumption, the polynomial given by Theorem 6 is of the form $(\det B(t)_{\{2,3\},\{2,3\}})g(t,x)$. The entries of $\operatorname{Sink}(B(t))$ are continuous functions of the entries of $B(t)$, and the roots of a polynomial are continuous functions of its coefficients, so $g(x)\mathrel{\mathop{:}}=\lim_{t\to 0^{+}}g(t,x)$ is a polynomial for the top-left entry of $\operatorname{Sink}(A)$. Moreover, since only one of the two cubic factors of $g(x)$ is independent of $s$, the top-left entry of $\operatorname{Sink}(A)$ is a root of that cubic factor. ∎
@@ 3. Square matrices and solving for coefficients
In this section, we use Theorem 6 to infer the form of an equation satisfied by the top-left entry of the Sinkhorn limit of a positive $n\times n$ matrix. We then interpolate the coefficients in the equation for $4\times 4$ matrices to obtain Conjecture 12.
For a positive $2\times 2$ matrix $A$, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies Equation (1), namely $\left(a_{11}a_{22}-a_{12}a_{21}\right)x^{2}-2a_{11}a_{22}\,x+a_{11}a_{22}=0$. Multiplying by $a_{11}$, this can be written using $\Delta$ and $\Gamma$ as
@@
Let $n\geq 1$. There exist integers $c_{S}(n)$, indexed by subsets $S\subseteq D(n,n)$, such that, for every positive $n\times n$ matrix $A$, the top-left entry $x$ of $\operatorname{Sink}(A)$ satisfies
@@
This conjecture implies that, since both $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ and $\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ are homogeneous degree-$(\lvert{R}\rvert+1)$ polynomials in the entries of $A$, each monomial $M(S)$ (and therefore also the coefficient of each $x^{k}$ in Equation (8)) is a homogeneous polynomial with degree
$\sum_{(R,C)\in D(n,n)}(\lvert{R}\rvert+1)=\sum_{k=0}^{n-1}\binom{n-1}{k}^{2}(k+1)=\frac{n+1}{2}\binom{2n-2}{n-1}.$
For $n=1,2,3,\dots$, this degree is $1,3,12,50,210,882,3696,15444,\dots$ [22, A092443].
The coefficients $c_{S}(n)$ in Conjecture 9 are not uniquely determined, since we can scale the polynomial. To remove this source of non-uniqueness, we define $c_{\{\}}(n)=1$ for all $n\geq 1$. Based on the polynomials for $n=2$ and $n=3$, we conjecture that $c_{D(n,n)}(n)=1$ for all $n\geq 2$.
We mention two surprising properties of the polynomial in Theorem 6. These properties also hold for the quadratic polynomial in Equation (6), so we expect them to hold more generally for the polynomial in Conjecture 9. The first is a symmetry in the coefficients. Each coefficient $d_{k}$ in Theorem 6 can be obtained from $d_{6-k}$ by replacing
@@
Conjecture 10.
Let $n\geq 2$. For each $(R,C)\in D(n,n)$, define $\kappa\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)=\begin{smallmatrix}\{2,3,\dots,n\}\setminus R\\ \{2,3,\dots,n\}\setminus C\end{smallmatrix}$, and extend $\kappa$ to subsets of $D(n,n)$ by
$\kappa\!\left(\!\begin{smallmatrix}R_{1}&R_{2}&\cdots&R_{k}\vphantom{\{}\\ C_{1}&C_{2}&\cdots&C_{k}\vphantom{\{}\end{smallmatrix}\!\right)=\left\{\kappa\!\left(\!\begin{smallmatrix}R_{1}\vphantom{\{}\\ C_{1}\vphantom{\{}\end{smallmatrix}\!\right),\kappa\!\left(\!\begin{smallmatrix}R_{2}\vphantom{\{}\\ C_{2}\vphantom{\{}\end{smallmatrix}\!\right),\dots,\kappa\!\left(\!\begin{smallmatrix}R_{k}\vphantom{\{}\\ C_{k}\vphantom{\{}\end{smallmatrix}\!\right)\right\}.$
For each $S\subseteq D(n,n)$, we have $c_{S}(n)=c_{\kappa(D(n,n)\setminus S)}(n)$.
The second is that, for each $k$, the sum of the coefficients $c_{S}(n)$ for $\lvert{S}\rvert=k$ is a signed binomial coefficient. For example, the coefficients in the linear combination (4) for $d_{1}$ sum to $-3-1-1-1-1+1=-6=-\binom{6}{1}$.
Conjecture 11.
For each $n\geq 1$, we have
$\sum_{S\subseteq D(n,n)}c_{S}(n)x^{\lvert{S}\rvert}=(x-1)^{\binom{2n-2}{n-1}}.$
@@
It remains to determine the coefficients $c_{S}(n)$. For $n=3$, we determined the coefficients $c_{S}(3)$ from the output of a Gröbner basis computation, but for $n\geq 4$ this computation seems to be infeasible. For a general $4\times 4$ matrix, we aborted the computation after $1$ week.
Instead, we generate many pseudorandom $n\times n$ matrices $A$, identify the top-left entry of $\operatorname{Sink}(A)$ as an algebraic number for each, and set up systems of linear equations in $c_{S}(n)$. We describe these three steps next.
@@
In the second step, for each matrix $A$ generated in the first step, we determine a polynomial equation satisfied by the top-left entry $x$ of $\operatorname{Sink}(A)$. There are two possible methods. One method is to use Gröbner bases; this is faster than the Gröbner basis computation for a matrix with symbolic entries, but for $n\geq 5$ it is still slow. Therefore we use another method, which is to apply the iterative scaling process to obtain a numeric approximation to $\operatorname{Sink}(A)$ and then guess a polynomial for its top-left entry. We begin by numericizing the integer entries of $A$ to high precision. For $n=4$ we use precision $2^{12}$, and for $n=5$ we use precision $2^{15}$. Then we iteratively scale until we reach a fixed point. The precision of the entries drops during the scaling process, but with the initial precisions $2^{12}$ and $2^{15}$ we get entries with sufficiently high precision that we can reliably recognize them using PSLQ. The expected degree of the polynomial is $d=\binom{2n-2}{n-1}$ according to Conjecture 9. Building in redundancy, we use Mathematica’s RootApproximant [24] to approximate $x$ by an algebraic number with target degree $d+2$. When the output has degree $d$, which is almost always the case, this is strong evidence that the approximation is in fact the exact algebraic number we seek. Occasionally the output has degree less than $d$, in which case we discard the matrix; for example, one $4\times 4$ matrix produced a polynomial with degree $8$ rather than $20$, presumably because the general degree-$20$ polynomial, when evaluated at the entries of the matrix, is reducible. Finally, we perform a check on the output by computing the ratio of its leading coefficient to its constant coefficient. This ratio should be $\frac{M(D(n,n))}{M(\{\})}$, assuming Conjecture 9 is correct and $c_{D(n,n)}(n)=1$. All outputs passed this test. We record the matrix $A$ along with the polynomial equation satisfied by $x$, and this will give us $1$ equation in the third step. (In fact we can get $n^{2}$ equations by applying PSLQ to each entry of the numeric approximation to $\operatorname{Sink}(A)$ and recording each polynomial along with the matrix obtained by swapping the appropriate rows and columns of $A$. For $n=5$ this is worthwhile, since our implementation of the iterative scaling process takes roughly $3$ minutes to reach a fixed point with initial precision $2^{15}$.)
@@
The third step is to determine the coefficients $c_{S}(n)$ in the coefficient of $x^{k}$ in Conjecture 9. For a given $k$, we do this by solving a system of linear equations involving the coefficients $c_{S}(n)$ where $\lvert{S}\rvert=k$. We assume $c_{T}(n)=c_{S}(n)$ if $T\equiv S$, so it suffices to determine $c_{S}(n)$ for one representative $S$ of each equivalence class, analogous to Theorem 6. This reduces the number of unknown coefficients, which reduces the number of equations we need, which reduces the number of matrices $A$ we apply the iterative scaling process to in the second step above. However, first we must partition $\{S\subseteq D(n,n):\lvert{S}\rvert=k\}$ into its equivalence classes under $\equiv$. Some care must be taken to do this efficiently; we make use of the fact that row permutations commute with column permutations, so it suffices to apply row permutations first. Once we have computed the equivalence classes, we take each polynomial computed in the second step above, scale it by an integer so that its leading coefficient is $M(D(n,n))$ (and therefore its constant coefficient is $M(\{\})$), extract the coefficient of $x^{k}$, and set this coefficient equal to $\sum_{S}c_{S}(n)\Sigma(S)$ where the sum is over one representative from each equivalence class of size-$k$ subsets. The number of equivalence classes tells us how many such equations we need in order to solve for the unknown coefficients $c_{S}(n)$. We include more equations than necessary in the system, building in redundancy, so that if a solution is found then we can be confident that the conjectured form is correct. Then we solve the system. In practice, even setting up the system can be computationally expensive. For $n=5$ and $k=4$, our initial implementation took $6$ days to set up the system (before solving!) because there are $1518$ unknown coefficients $c_{S}(n)$, so we need at least that many equations, and each equation involves $\binom{70}{4}=916895$ monomials, each of which is a product of $\binom{8}{4}=70$ determinants. For small $k$, a more efficient way to construct each equation is to take advantage of the fact that, for each pair $S,T$ of subsets of $D(5,5)$, the products $M(S)$ and $M(T)$ have most factors in common, and almost all are $\Gamma$ factors. Therefore, we can reduce the number of determinant computations by dividing both sides by $M(\{\})=\prod_{(R,C)\in D(5,5)}\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$, which is nonzero by the assumption that no minor is 00. On the left, we compute $M(\{\})$ once and divide the extracted coefficient of $x^{k}$ by $M(\{\})$. On the right, we divide each $M(S)$ by $M(\{\})$. Instead of computing $M(S)/M(\{\})$ from definitions, we precompute the ratio $\Delta\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)/\Gamma\!\left(\!\begin{smallmatrix}R\vphantom{\{}\\ C\vphantom{\{}\end{smallmatrix}\!\right)$ for each $(R,C)\in D(5,5)$; then, for each $S$ with $\lvert{S}\rvert=k$, we use the precomputed ratios to compute
@@
This product can be computed quickly small since $k$ is small. (This shortcut is also used by SinkhornPolynomials [19] to compute polynomials more quickly when no minor is 00.)
We now carry out these three steps for $n=4$ to obtain a polynomial for the top-left entry $x$ of $\operatorname{Sink}(A)$ for $4\times 4$ matrices $A$. This polynomial is analogous to the polynomial in Theorem 6 for $3\times 3$ matrices. The number of equivalence classes of size-$k$ subsets of $D(4,4)$ for $k=0,1,\dots,20$ is
@@
The number of unknown coefficients $c_{S}(4)$ is the sum of these numbers, which is $17920$. We use the definition $c_{\{\}}(4)=1$ and the conjecture $c_{D(4,4)}(4)=1$, and we solve the remaining $19$ systems of linear equations, the largest of which requires $3024$ equations. Unfortunately, for each $k\in\{4,5,\dots,16\}$ the system has multiple solutions. For example, when we solve the system for $k=4$, only $104$ of the $123$ unknown coefficients $c_{S}(4)$ with $\lvert{S}\rvert=4$ are uniquely determined; the other $19$ are parameterized by $2$ free variables. For $k=10$, the solution space has dimension $1141$. This is not a weakness of the interpolation strategy but rather implies that the coefficients of $x^{4},x^{5},\dots,x^{16}$ have multiple representations as linear combinations of class sums $\Sigma(S)$ and therefore that relations exist among these class sums. However, by setting the free variables to 00 (or any other values), we obtain numeric coefficients $c_{S}(4)$.
@@
such that the top-left entry $x$ of the Sinkhorn limit of a $4\times 4$ matrix satisfies
$\sum_{k=0}^{20}\left(\sum_{S}c_{S}(4)\Sigma(S)\right)x^{k}=0$
@@
and let $x$ be the top-left entry of $\operatorname{Sink}(A)$. Conjecture 12 gives
@@
In principle, we can use the same method to interpolate a polynomial equation for $n\times n$ matrices for any given $n$. However, the computation is formidable. For $5\times 5$ matrices, we were able to compute the coefficients for $k=4$ and $k=5$ (where the families of representations are respectively $8$-dimensional and $44$-dimensional), but for $k=6$ we could not get the computation of equivalence classes to finish (aborted after $2$ weeks). Completing the computation up to the halfway point $k=\frac{1}{2}\binom{8}{4}=35$, after which we could use the symmetry in Conjecture 10, seems to be infeasible.